# PAC Apprenticeship Learning with Bayesian Active Inverse Reinforcement Learning

### v1

@@ Abstract
As AI systems become increasingly autonomous, reliably aligning their decision-making to human preferences is essential. Inverse reinforcement learning (IRL) offers a promising approach to infer preferences from demonstrations. These preferences can then be used to produce an apprentice policy that performs well on the demonstrated task. However, in domains like autonomous driving or robotics, where errors can have serious consequences, we need not just good average performance but reliable policies with formal guarantees – yet obtaining sufficient human demonstrations for reliability guarantees can be costly. Active IRL addresses this challenge by strategically selecting the most informative scenarios for human demonstration. We introduce PAC-EIG, an information-theoretic acquisition function that directly targets probably-approximately-correct (PAC) guarantees for the learned policy – providing the first such theoretical guarantee for active IRL with noisy expert demonstrations. Our method maximises information gain about the regret of the apprentice policy, efficiently identifying states requiring further demonstration. We also present Reward-EIG as an alternative when learning the reward itself is the primary objective. Focusing on finite state-action spaces, we prove convergence bounds, illustrate failure modes of prior heuristic methods, and demonstrate our method’s advantages experimentally.
@@ 1 Introduction
Stuart Russell suggested three principles for the development of beneficial artificial intelligence: its only objective is realizing human preferences, it is initially uncertain about these preferences, and its ultimate source of information about them is human behavior (Russell, 2019). Apprenticeship learning via Bayesian inverse reinforcement learning (IRL) can be understood as a possible operationalization of these principles: Bayesian IRL starts with a prior distribution over reward functions representing initial uncertainty about human preferences. It then combines this prior with demonstration data from a human expert acting approximately optimally with respect to the unknown reward, to produce a posterior distribution over rewards. In apprenticeship learning, this posterior over rewards is then used to produce a policy that should perform well with respect to the unknown reward function.
@@ 2 Task formulation
Let $\mathcal{M}=\left(\mathcal{S},\mathcal{A},p,r,\gamma,t_{\max},\rho\right)$ be a parameterized Markov decision process (MDP), where $\mathcal{S}$ and $\mathcal{A}$ are finite state and action spaces respectively, $p:\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is the transition function where $\mathcal{P}(\mathcal{S})$ is a set of probability measures over $\mathcal{S}$, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is an (expected) reward function, [^1] Our formulation permits the reward to be stochastic. However, our expert model (1) depends on the rewards only via the optimal Q-function, which in turn depends only on the expected reward. Thus, the demonstrations can only ever give us information about the expectation. Throughout the paper, the learnt reward function can be interpreted either as modeling a deterministic reward, or an expectation of a stochastic reward. $\gamma\in(0,1)$ is a discount rate, $t_{\max}\in\mathbb{N}\cup\{\infty\}$ is the time horizon, and $\rho$ is the initial state distribution. We assume the learner has full knowledge of the MDP except for the reward.
@@
These objectives are often closely connected – learning about the reward function enables improving the apprentice policy. However, espcially in the active setting, they can come apart – for instance, once we know an action $a$ leads to lower return than $a^{\prime}$ in a particular state, we may no longer need to gather further information about rewards in these states for the apprenticeship learning objective as we already know to choose the better action, while the reward-learning objective may motivate further queries to reduce the reward uncertainty.
@@
By $V_{r}^{\pi}$ we denote the state-value function of policy $\pi$ with respect to reward $r$. $V_{r}^{*}$ is then the value function of the optimal policy with respect to $r$. A lack of subscript, as in $V^{*}$, indicates (optimal) value with respect to the true reward (since the true reward is not known by the learner, this generally needs to be treated as random variable). By $G_{r}(\tau)$ we denote the return of trajectory $\tau$ with respect to $r$. By $R^{\pi}_{r}(s_{0})$ we denote the regret of policy $\pi$ starting from state $s_{0}$, i.e. $R^{\pi}_{r}(s_{0}):=V^{*}_{r}(s_{0})-V^{\pi}_{r}(s_{0})$, and $R^{\pi}_{r}:=\mathbb{E}_{s_{0}\sim\rho}R^{\pi}_{r}(s_{0})$. We also call immediate regret the quantity $R^{*}_{\pi,r}(s)=V^{*}(s)-Q^{*}(s,\pi(s))$ and also denote by $R^{*}_{\pi,r}(s,a)=\max\{0,Q^{*}(s,a)-Q^{*}(s,\pi(s))\}$ the immediate regret relative to action $a$ in state $s$. You can also find a table with the notation used in this paper in Appendix F.
@@ 4.1 PAC-EIG: Information Gain for Reliable Policies
Our goal is to produce an apprentice policy that is probably approximately correct – that is, with high probability ($1-\delta$), the policy’s regret is bounded by $\epsilon$. To achieve this efficiently, we need to identify states where the current apprentice policy might be making poor decisions. To this end, we define the immediate regret of an apprentice policy $\pi^{A}$ in state $s$ as:
$R^{*}_{\pi^{A},r}(s)=V^{*}_{r}(s)-Q^{*}_{r}(s,\pi^{A}(s))$ (4)
which captures how much value we lose by following the apprentice policy in state $s$ compared to optimal behavior. This can be decomposed per action as $R^{*}_{\pi^{A},r}(s,a)=\max\{0,Q^{*}_{r}(s,a)-Q^{*}_{r}(s,\pi^{A}(s))\}$, representing the regret relative to choosing a particular alternative action $a$. This regret is unknown to us, so we need to treat it as a random variable.
For computational tractability, we discretize the immediate regret into a ternary variable $E_{s,a}$ tracking state- and action-wise correctness as follows:
$E_{s,a}^{{\pi^{\text{A}}}}=\begin{cases}\text{``correct''}&\text{if }R^{*}_{\pi^{A},r}(s,a)=0\\ \text{``approximately correct''}&\text{if }0<R^{*}_{\pi^{A},r}(s,a)<\epsilon(1-\gamma)\\ \text{``not correct''}&\text{if }R^{*}_{\pi^{A},r}(s,a)\geq\epsilon(1-\gamma)\end{cases}$ (5)
@@
While we are aiming for a $\epsilon$-$\delta$-PAC policy as the final output, for intermediate steps, our algorithm uses an apprentice policy ${\pi^{\text{A}}}(s)=\text{argmax}_{a}\mathbb{P}[Q^{*}(s,\pi^{*}(s))-Q^{*}(s,a)=0|\mathcal{D}_{n}],\;\forall s\in\mathcal{S}$, i.e. an apprentice policy maximizes the probability of taking the optimal action in each step. The reason for this choice is that it ensures there is probability mass on both of two contrastive options: the policy being optimal in any given state on the one hand, and, on the other, it having an immediate regret of at least $(1-\gamma)\epsilon$ in at least one state as long as the PAC condition is not satisfied. This contrast ensures that we gain information by observing the expert in this state as we detail in the next section.
Especially if we care about gradually tightening the bound (rather than just reaching a particular bound) we may wish to dynamically set $\epsilon$ in each step to just below the threshold for which the condition is currently satisfied.
@@ 4.2 Computing PAC-EIG
    For each Q-value sample, compute the discretized regret values $E^{{\pi^{\text{A}}}}_{s,a}$ for all state-action pairs. Note that multiple Q-value samples may map to the same discretized configuration $E^{{\pi^{\text{A}}}}$, so there are at most $M_{E}\leq M$ distinct values of $E^{{\pi^{\text{A}}}}$.
@@
    Estimate the expected information gain using the standard Monte Carlo estimator as
@@
    where the trajectory probability given $E^{{\pi^{\text{A}}}}$ can be computed as $p(\tau|E^{{\pi^{\text{A}}}},s_{0})=\prod_{(s_{t},a_{t})\in\tau}p(a_{t}|s_{t},E)$ (omitting the transition probabilities since they would cancel out in the log-ratio). To compute $p(a_{t}|s_{t},E^{{\pi^{\text{A}}}})$, we average the expert action probabilities over all Q-value samples that map to the same discretized configuration: $p(a_{t}|s_{t},E^{{\pi^{\text{A}}}})=\frac{1}{|\mathcal{Q}_{E}^{{\pi^{\text{A}}}}|}\sum_{Q^{*}\in\mathcal{Q}_{E}}p(a_{t}|s_{t},Q^{*})$, where $\mathcal{Q}_{E}$ denotes the set of Q-value samples corresponding to the discretized regret configuration $E^{{\pi^{\text{A}}}}$.
See Appendix C for additional implementation details.
@@ 4.3 Reward EIG: When Learning the Reward is the Goal
While our primary focus above has been on producing reliable apprentice policies, in some applications the reward function itself is of intrinsic interest – for instance, when using IRL to understand animal behavior (Ashwood et al., 2022) or human preferences (Huang et al., 2022). For these cases, we can still use the EIG framework, but instead maximize the expected information gain about the reward:
@@
See Appendix D for a detailed discussion of alternative acquisition functions and the theoretical connections between different formulations.
@@ 5 Producing a PAC Policy
For $\epsilon>0$ and $\delta\in(0,\frac{1}{2}]$, assume that no policy $\pi$ is $(\epsilon,\delta)$-probably-approximately-correct, i.e., $\mathbb{P}[R^{\pi}_{r}\geq\epsilon]>\delta,\;\forall\pi$. Then, there exists a state $s\in\mathcal{S}$ such that observing a new expert demonstration at $s$ has an expected information gain with respect to the variable $E^{{\pi^{\text{A}}}}$ (Eq. 5) of at least
@@
For the ternary discretization used in PAC-EIG, the maximum initial entropy is $\log(3)|\mathcal{S}||\mathcal{A}|$ for any given policy and the associated variable $E^{{\pi^{\text{A}}}}$. However, since the policy may change during the optimization process, we get a total initial entropy across all possible policies of $h_{\text{max}}\leq\log(3)|\mathcal{S}||\mathcal{A}|^{2}$ (even though there are $|\mathcal{S}||\mathcal{A}|$ possible policies, note that the ternary variable in a given state does not depend on the policy’s actions in other states, which allows us to keep the $|\mathcal{S}|$ term instead of a $|\mathcal{S}|^{2}$).
@@ 6 Experiments
Each experiment type was run with with 16 different random reward functions, different terminal states (except for the jail environment) and different 2 initial states in the 10x10 environment. The plots display the mean and the standard error across these 16 random instances.
@@ 6.1 Results
Results on the simple environment from Figure 1 illustrate a crucial failure mode of the action entropy (Kweon et al., 2023) acquisition function – it always queries the jail state and thus fails to learn anything useful, while both reward and regret EIG learn an optimal policy within 10 steps in all 16 instances with similar posterior entropies.
@@
In the case of only 2 initial states on the 10x10 gridworld, we can see our regret-focused PAC-EIG acquisition function to do much better in terms of both actual and posterior regret, reaching a zero true regret, as well as 0.1-0.1-PAC apprentice policy, by step 50. ActiveVaR remains competitive with RewardEIG in terms of the reward-learning objective (entropy of the rewardposterior), but both fall behind PAC-EIG in terms of true and posterior regret.
On the gridworld with fully uniform initial states, we observe that both our information theoretic acquisition functions result in lower posterior reward entropy and lower regret than prior methods except for ActiveVaR, which seems to roughly match their performance. Interestingly, the reward-based acquisition function and the regret-based ones seem to perform similarly well on both objectives, suggesting that there is a strong correlation between learning about the reward and learning about the apprentice regret in this environment with uniform initial states.
@@ 7 Discussion and conclusion
In this paper we have proposed new acquisition functions for active IRL, each geared toward one of two possible objectives: learning about an unknown reward function, or producing a well-performing apprentice policy. We have shown that across a set of gridworld experiments, our acquisition functions outperform or at least match prior methods on their respective objective. Furthermore, our immediate-regret EIG acquisition function is a first acquisition function with a regret bound in our setting. While we have so far tested the methods only in finite state spaces, both of them were constructed to generalize also to continuous spaces, which will be addressed in future work.
@@ Appendix B Theoretical Analysis
We will establish an upper bound on the expected number of expert demonstrations needed to find a policy satisfying an $\epsilon$-$\delta$-PAC criterion using the PAC-EIG acquisition function. As outlined in Section 4, we will assume that during the learning process, the apprentice policy is one maximizing the probability of being $\epsilon(1-\gamma)/2$-approximately correct in each state in the immediate regret sense, i.e.
${\pi^{\text{A}}}(s)=\text{argmax}_{a}\mathbb{P}[Q*(s,\pi^{*}(s))-Q^{*}(s,a)<\epsilon(1-\gamma)/2|\mathcal{D}_{n}],\;\forall s\in\mathcal{S}$ (18)
. The proof strategy proceeds in three steps:
@@
    Building on this, we prove that if a policy is not $(\epsilon,\delta)$-PAC, then there exists a state where the difference between optimal and apprentice policy’s optimal Q-values is lower-bounded by a function of $\epsilon$ with probability at least $\delta/|\mathcal{S}|$.
@@
    Finally, we show that in such cases, observing an expert demonstration from an appropriately chosen initial state provides a guaranteed minimum amount of information about whether the policy is approximately correct. Since we can only gain a finite amount of information (bounded by the entropy of our prior), this leads to a bound on the number of demonstrations needed.
We begin with our first lemma, which connects overall policy regret to statewise immediate regret, i.e. differences in optimal Q-values:
@@
($e_{\text{n}}$ standing for the appsince the action since $a^{\prime}$ being optimal is a subset of the event $e_{\text{n}}$ conditioned on $e_{*}$.
@@
For completeness, we can denote by $e_{\text{a}}$ the "approximately correct" complement of the events $e_{\text{c}}$ and $e_{\text{o}}$.
@@
$I(A;E_{s,a^{\prime}})$ $=\sum_{e\in\{e_{\text{c}},e_{\text{a}},e_{\text{n}}\}}\mathbb{P}[e]D_{\text{KL}}(p(A|e)\|p(A))$
@@
If $a^{\prime}$ is optimal, its probability of being selected by the expert is at least $1/|\mathcal{A}|$. Since $a^{\prime}$ was chosen so that it has a probability of at least $\frac{1}{|\mathcal{A}|-1}$ of being optimal conditional on $e_{*}$, it must also have a probability of at least $\frac{1}{|\mathcal{A}|-1}$ conditional on $e_{\text{n}}$ (since $e_{\text{n}}\subseteq e_{*}$ and $\{a\;\text{optimal}\}\cap e_{*}\subseteq e_{n}$). Thus, we have
@@
$\geq\frac{1}{|\mathcal{A}|(|\mathcal{A}|-1)}.$
@@
$\mathbb{P}\!\left[A=a^{\prime}|e_{\text{n}}\right]-\mathbb{P}\!\left[A={\pi^{\text{A}}}(s)|e_{\text{n}}\right]$ $>\frac{1-e^{-\beta(1-\gamma)\epsilon}}{|\mathcal{A}|(|\mathcal{A}|-1)}$
@@
$D_{\text{TV}}(p(A|e_{\text{c}}),p(A|e_{\text{n}}))$ $\geq\frac{1}{2}\left(\left|\mathbb{P}\left[A=a^{\prime}|e_{\text{n}}\right]+\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{n}}\right]\right|-\left|\mathbb{P}\left[A=a^{\prime}|e_{\text{c}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{c}}\right]\right|\right)$
$\geq\frac{1}{2}\left(\mathbb{P}\left[A=a^{\prime}|e_{\text{n}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{n}}\right]\right)-\left(\mathbb{P}\left[A=a^{\prime}|e_{\text{c}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{c}}\right]\right)$
@@
where we used the definition of total variation distance in the first step (omitting non-negative terms corresponding to actions other than ${\pi^{\text{A}}}(s)$ and $a^{\prime}$), the triangle inequality and rearranging in the second step, and, in the last step, using the lower bound we just derived and dropping the second term, since $\mathbb{P}\left[A=a^{\prime}|e_{\text{c}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{c}}\right]\leq 0$.
@@
$\geq D_{\text{TV}}(p(A|e_{\text{n}}),p(A|e_{\text{c}})))^{2}$ $>\frac{1}{4|\mathcal{A}|^{4}}e^{-\beta(1-\gamma)\epsilon}(1-e^{-\beta(1-\gamma)\epsilon/2})^{2}.$
where we applied the inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$ in the second step, and the triangle inequality in the fourth.
@@
$I(A;E_{s,a^{\prime}})$ $=\sum_{E_{s,a^{\prime}}\in\{e_{\text{n}},e_{\text{c}},e_{\text{w}}\}}\mathbb{P}[E_{s,a^{\prime}}]D_{\text{KL}}(p(A|E_{s,a^{\prime}})\|p(A))$

### v2

@@ Abstract
As AI systems become increasingly autonomous, reliably aligning their decision-making with human preferences is essential. Inverse reinforcement learning (IRL) offers a promising approach to infer preferences from demonstrations. These preferences can then be used to produce an apprentice policy that performs well on the demonstrated task. However, in domains like autonomous driving or robotics, where errors can have serious consequences, we need not just good average performance but reliable policies with formal guarantees – yet obtaining sufficient human demonstrations for reliability guarantees can be costly. Active IRL addresses this challenge by strategically selecting the most informative scenarios for human demonstration. We introduce PAC-EIG, an information-theoretic acquisition function that directly targets probably-approximately-correct (PAC) guarantees for the learned policy – providing the first such theoretical guarantee for active IRL with noisy expert demonstrations. Our method maximises information gain about the regret of the apprentice policy, efficiently identifying states requiring further demonstration. We also present Reward-EIG as an alternative when learning the reward itself is the primary objective. Focusing on finite state-action spaces, we prove convergence bounds, illustrate failure modes of prior heuristic methods, and demonstrate our method’s advantages experimentally.
@@ 1 Introduction
Stuart Russell suggested three principles for the development of beneficial artificial intelligence: its only objective is to realize human preferences, it is initially uncertain about these preferences, and its ultimate source of information about them is human behaviour (Russell, 2019). Apprenticeship learning via Bayesian inverse reinforcement learning (IRL) can be understood as a possible operationalization of these principles: Bayesian IRL starts with a prior distribution over reward functions representing initial uncertainty about human preferences. It then combines this prior with demonstration data from a human expert acting approximately optimally with respect to the unknown reward, to produce a posterior distribution over rewards. In apprenticeship learning, this posterior over rewards is then used to produce a policy that should perform well with respect to the unknown reward function.
@@ 2 Task formulation
Let $\mathcal{M}=\left(\mathcal{S},\mathcal{A},p,r,\gamma,t_{\max},\rho\right)$ be a parameterized Markov decision process (MDP), where $\mathcal{S}$ and $\mathcal{A}$ are finite state and action spaces respectively, $p:\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is the transition function where $\mathcal{P}(\mathcal{S})$ is a set of probability measures over $\mathcal{S}$, $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is an (expected) reward function, [^1] Our formulation permits the reward to be stochastic. However, our expert model (1) depends on the rewards only via the optimal Q-function, which in turn depends only on the expected reward. Thus, the demonstrations can only ever give us information about the expectation. Throughout the paper, the learnt reward function can be interpreted either as modelling a deterministic reward, or an expectation of a stochastic reward. $\gamma\in(0,1)$ is a discount rate, $t_{\max}\in\mathbb{N}\cup\{\infty\}$ is the time horizon, and $\rho$ is the initial state distribution. We assume the learner has full knowledge of the MDP except for the reward.
@@
These objectives are often closely connected – learning about the reward function enables improving the apprentice policy. However, especially in the active setting, they can come apart – for instance, once we know an action $a$ leads to lower return than $a^{\prime}$ in a particular state, we may no longer need to gather further information about rewards in these states for the apprenticeship learning objective as we already know to choose the better action, while the reward-learning objective may motivate further queries to reduce the reward uncertainty.
@@
By $V_{r}^{\pi}$ we denote the state-value function of policy $\pi$ with respect to reward $r$. $V_{r}^{*}$ is then the value function of the optimal policy with respect to $r$. A lack of subscript, as in $V^{*}$, indicates (optimal) value with respect to the true reward (since the true reward is not known by the learner, this generally needs to be treated as random variable). By $G_{r}(\tau)$ we denote the return of trajectory $\tau$ with respect to $r$. By $R^{\pi}_{r}(s_{0})$ we denote the regret of policy $\pi$ starting from state $s_{0}$, i.e. $R^{\pi}_{r}(s_{0}):=V^{*}_{r}(s_{0})-V^{\pi}_{r}(s_{0})$, and $R^{\pi}_{r}:=\mathbb{E}_{s_{0}\sim\rho}R^{\pi}_{r}(s_{0})$. We also call immediate regret the quantity $R^{*}_{\pi,r}(s)=V^{*}(s)-Q^{*}(s,\pi(s))$ and also denote by $R^{*}_{\pi,r}(s,a)=\max\{0,Q^{*}(s,a)-Q^{*}(s,\pi(s))\}$ the immediate regret relative to action $a$ in state $s$. You can also find a table with the notation used in this paper in Appendix E.
@@ 4.1 PAC-EIG: Information Gain for Reliable Policies
Our goal is to produce an apprentice policy that is probably approximately correct – that is, with high probability ($1-\delta$), the policy’s regret is bounded by $\epsilon$. To achieve this efficiently, we need to identify states where the current apprentice policy might be making poor decisions. To this end, we define the immediate regret of an apprentice policy $\pi^{\text{A}}$} in state $s$ as:
$R^{*}_{{\pi^{\text{A}}},r}(s)=V^{*}_{r}(s)-Q^{*}_{r}(s,{\pi^{\text{A}}}(s))$ (4)
which captures how much value we lose by following the apprentice policy in state $s$ compared to optimal behaviour. This can be decomposed per action as
$R^{*}_{{\pi^{\text{A}}},r}(s,a)=\max\{0,Q^{*}_{r}(s,a)-Q^{*}_{r}(s,{\pi^{\text{A}}}(s))\},$
representing the regret relative to choosing a particular alternative action $a$. This regret is unknown to us, so we need to treat it as a random variable. We propose to use the expected information gain about this immediate regret as a theoretical acquisition function that helps us find a PAC policy.
For practical computation, we discretize the immediate regret into a ternary variable $E_{s,a}$ tracking state- and action-wise correctness as follows:
$E_{s,a}^{{\pi^{\text{A}}}}=\begin{cases}\text{``correct''}&\text{if }R^{*}_{{\pi^{\text{A}}},r}(s,a)=0\\ \text{``approximately correct''}&\text{if }0<R^{*}_{{\pi^{\text{A}}},r}(s,a)<\epsilon(1-\gamma)\\ \text{``not correct''}&\text{if }R^{*}_{{\pi^{\text{A}}},r}(s,a)\geq\epsilon(1-\gamma)\end{cases}$ (5)
@@
While we are aiming for a $\epsilon$-$\delta$-PAC policy as the final output, for intermediate steps, our algorithm uses an apprentice policy ${\pi^{\text{A}}}(s)=\text{argmax}_{a}\mathbb{P}[Q^{*}(s,\pi^{*}(s))-Q^{*}(s,a)=0|\mathcal{D}_{n}],\;\forall s\in\mathcal{S}$, i.e. an apprentice policy maximizes the probability of taking the optimal action in each state. The reason for this choice is that it ensures there is probability mass on both of two contrastive options: the policy being optimal in any given state on the one hand, and, on the other, it having an immediate regret of at least $(1-\gamma)\epsilon$ in at least one state as long as the PAC condition is not satisfied. This contrast ensures that we gain information by observing the expert in this state as we detail in the next section.
[…]
@@ 4.2 Computing PAC-EIG
    For each Q-value sample, compute the discretized regret values $E^{{\pi^{\text{A}}}}$. Note that multiple Q-value samples may map to the same discretized configuration $E^{{\pi^{\text{A}}}}$, so there are at most $M_{E}\leq M$ distinct values for the samples of $E^{{\pi^{\text{A}}}}$.
@@
    Estimate the expected information gain using the standard Monte Carlo estimator of EIG as
@@
    where the trajectory probability given $E^{{\pi^{\text{A}}}}$ can be computed as $p(\tau|E^{{\pi^{\text{A}}}},s_{0})=\prod_{(s_{t},a_{t})\in\tau}p(a_{t}|s_{t},E^{\pi^{\text{A}}})$ (omitting the transition probabilities since they would cancel out in the log-ratio). To compute $p(a_{t}|s_{t},E^{{\pi^{\text{A}}}})$, we average the expert action probabilities over all Q-value samples that map to the same discretized configuration: $p(a_{t}|s_{t},E^{{\pi^{\text{A}}}})=\frac{1}{|\mathcal{Q}_{E^{\pi^{\text{A}}}}|}\sum_{Q^{*}\in\mathcal{Q}_{E^{\pi^{\text{A}}}}}p(a_{t}|s_{t},Q^{*})$, where $\mathcal{Q}_{E^{\pi^{\text{A}}}}$ denotes the set of Q-value samples corresponding to the discretized regret configuration $E^{{\pi^{\text{A}}}}$ and the action probability the Boltzmann rational expert policy 1.
@@ 4.3 Reward EIG: When Learning the Reward is the Goal
While our primary focus above has been on producing reliable apprentice policies, in some applications the reward function itself is of intrinsic interest – for instance, when using IRL to understand animal behaviour (Ashwood et al., 2022) or human preferences (Huang et al., 2022). For these cases, we can still use the EIG framework, but instead maximize the expected information gain about the reward:
@@ 5 Producing a PAC Policy
For $\epsilon>0$ and $\delta\in(0,\frac{1}{2}]$, assume that no policy $\pi^{\text{A}}$} is $(\epsilon,\delta)$-probably-approximately-correct, i.e., $\mathbb{P}[R^{\pi}_{r}\geq\epsilon]>\delta,\;\forall\pi$. Then, there exists a state $s\in\mathcal{S}$ such that observing a new expert demonstration at $s$ has an expected information gain with respect to the variable $E^{{\pi^{\text{A}}}}$ (Eq. 5) of at least
@@
For the ternary discretization used in PAC-EIG, the maximum initial entropy is $\log(3)|\mathcal{S}|(|\mathcal{A}|-1)$ for any given policy and the associated variable $E^{{\pi^{\text{A}}}}$. However, since the policy may change during the optimization process, we get a total initial entropy across all possible policies of $h_{\text{max}}\leq\log(3)|\mathcal{S}||\mathcal{A}|(|\mathcal{A}|-1)$ (even though there are $|\mathcal{A}|^{|\mathcal{S}|}$ possible policies, note that the ternary variable in a given state does not depend on the policy’s actions in other states, leaving the number of distinct components of $E$ at $|\mathcal{S}||\mathcal{A}|(|\mathcal{A}|-1)$).
@@ 6 Experiments
Each experiment type was run with 16 different random seeds (i.e. combinations of random reward functions, terminal states (except for the jail environment), and different 2 initial states in the 10x10 environment). The plots display the mean and the standard error across these 16 random instances.
@@ 6.1 Results
Figure 2 shows results on the simple environment from Figure 1 to illustrate a crucial failure mode of the action entropy (Kweon et al., 2023) acquisition function – it always queries the jail state and thus fails to learn anything useful, while both reward and regret EIG learn an optimal policy within 10 steps in all 16 instances with similar posterior entropies.
@@
[…]
On the 8x8 gridworld with fully uniform initial states, we observe that both our information-theoretic acquisition functions result in lower posterior reward entropy and lower regret than prior methods except for ActiveVaR, which seems to roughly match their performance. Interestingly, the reward-based acquisition function and the regret-based ones seem to perform similarly well on both objectives, suggesting that there is a strong correlation between learning about the reward and learning about the apprentice regret in this environment with uniform initial states.
@@ 7 Discussion and conclusion
In this paper, we have proposed new acquisition functions for active IRL, each geared toward one of two possible objectives: learning about an unknown reward function, or producing a well-performing apprentice policy. We have shown that across a set of gridworld experiments, our acquisition functions outperform or at least match prior methods on their respective objectives. Furthermore, our immediate-regret EIG acquisition function is the first acquisition function with a regret bound in our setting. While we have so far tested the methods only in finite state spaces, both of them were constructed to generalize also to continuous spaces, which will be addressed in future work.
@@ Appendix A Failure modes of prior methods
While this is a possible conceptual objection to this acquisition function (which could be addressed by moving from regret to our immediate regret), we think this acquisition function is a strong option, which is also confirmed by empirical results.
@@ Appendix B Theoretical Analysis
We will establish an upper bound on the expected number of expert demonstrations needed to find a policy satisfying an $\epsilon$-$\delta$-PAC criterion using the PAC-EIG acquisition function. As outlined in Section 4, we will assume that during the learning process, the apprentice policy is one that in each state, maximizes the probability of taking an optimal action. The proof strategy proceeds in three steps:
@@
    Building on this, we prove that if a policy is not $(\epsilon,\delta)$-PAC, then there exists a state where the difference between optimal and apprentice policy’s optimal Q-values is lower-bounded by $\epsilon(1-\gamma)$ with probability at least $\delta/|\mathcal{S}|$.
@@
    Finally, we show that in such cases, observing an expert demonstration from an appropriately chosen initial state provides a guaranteed minimum amount of information about whether the policy is approximately correct. Since wea function of can only gain a finite amount of information (bounded by the entropy of our prior), this leads to a bound on the number of demonstrations needed.
We begin with our first lemma, which connects overall policy regret to statewise immediate regret, i.e. differences in optimal Q-values. This could be viewed as a corollary of the Performance Difference Lemma Kakade & Langford (2002), but we will prove it directly for completeness.
@@
($e_{\text{n}}$ standing for the apprentice action not being approximately correct relative to $a^{\prime}$) since the action $a^{\prime}$ being optimal is a subset of the event $e_{\text{n}}$ conditioned on $e_{*}$.
@@
For completeness, we can denote by $e_{\text{a}}$ the "approximately correct" complement of the events $e_{\text{c}}$ and $e_{\text{n}}$.
@@
$I(A;E^{\pi^{\text{A}}}_{s,a^{\prime}})$ $=\sum_{e\in\{e_{\text{c}},e_{\text{a}},e_{\text{n}}\}}\mathbb{P}[e]D_{\text{KL}}(p(A|e)\|p(A))$
@@
If $a^{\prime}$ is optimal, its probability of being selected by the expert is at least $1/|\mathcal{A}|$. In fact, conditional on $e_{\text{n}}$, its probability must be strictly greater than that of ${\pi^{\text{A}}}(s)$, so its probability of being selected if optimal is strictly greater than $\frac{1}{|\mathcal{A}|}$. Since $a^{\prime}$ was chosen so that it has a probability of at least $\frac{1}{|\mathcal{A}|-1}$ of being optimal conditional on $e_{*}$, it must also have a probability of at least $\frac{1}{|\mathcal{A}|-1}$ conditional on $e_{\text{n}}$ (since $e_{\text{n}}\subseteq e_{*}$ and $\{a\;\text{optimal}\}\cap e_{*}\subseteq e_{n}$). Thus, we have
@@
$>\frac{1}{|\mathcal{A}|(|\mathcal{A}|-1)}.$
@@
$\mathbb{P}\!\left[A=a^{\prime}|e_{\text{n}}\right]-\mathbb{P}\!\left[A={\pi^{\text{A}}}(s)|e_{\text{n}}\right]$ $>\frac{1-e^{-\beta(1-\gamma)\epsilon}}{|\mathcal{A}|(|\mathcal{A}|-1)}\;.$
@@
$D_{\text{TV}}$ $\left(p(A|e_{\text{c}}),p(A|e_{\text{n}})\right)$
$\geq\frac{1}{2}\biggl(\Bigl|\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{c}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{n}}\right]\Bigr|+\Bigl|\mathbb{P}\left[A=a^{\prime}|e_{\text{n}}\right]-\mathbb{P}\left[A=a^{\prime}|e_{\text{c}}\right]\Bigr|\biggr)$
$\geq\frac{1}{2}\biggl(\Bigl(\mathbb{P}\left[A=a^{\prime}|e_{\text{n}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{n}}\right]\Bigr)-\Bigl(\mathbb{P}\left[A=a^{\prime}|e_{\text{c}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{c}}\right]\Bigr)\biggr)$
@@
where we used the definition of total variation distance in the first step (omitting non-negative terms corresponding to actions other than ${\pi^{\text{A}}}(s)$ and $a^{\prime}$), the triangle inequality, dropping absolute value, and rearranging terms in the second step, and, in the last step, using the lower bound we just derived and dropping the second term, since $\mathbb{P}\left[A=a^{\prime}|e_{\text{c}}\right]-\mathbb{P}\left[A={\pi^{\text{A}}}(s)|e_{\text{c}}\right]\leq 0$.
@@
$\geq D_{\text{TV}}(p(A|e_{\text{n}}),p(A|e_{\text{c}})))^{2}$
$>\frac{1}{4|\mathcal{A}|^{2}(|\mathcal{A}|-1)^{2}}(1-e^{-\beta(1-\gamma)\epsilon})^{2}.$
where we applied the inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$ in the second step, and the triangle inequality in the third.
@@
$I(A;E^{\pi^{\text{A}}}_{s,a^{\prime}})$ $=\sum_{E^{\pi^{\text{A}}}_{s,a^{\prime}}\in\{e_{\text{n}},e_{\text{c}},e_{\text{a}}\}}\mathbb{P}[E^{\pi^{\text{A}}}_{s,a^{\prime}}]D_{\text{KL}}(p(A|E^{\pi^{\text{A}}}_{s,a^{\prime}})\|p(A))$